# Auditory–Tactile Congruence for Synthesis of Adaptive Pain Expressions in RoboPatients

### v1

@@ Abstract
Misdiagnosis can lead to delayed treatments and harm. Robotic patients offer a controlled way to train and evaluate clinicians in rare, subtle, or complex cases, reducing diagnostic errors. We present RoboPatient, a medical robotic simulator aimed at multimodal pain synthesis based on haptic and auditory feedback during palpation-based training scenarios. The robopatient functions as an adaptive intermediary, capable of synthesizing plausible pain expressions—vocal and facial—in response to tactile stimuli generated during palpation. Using an abdominal phantom, robopatient captures and processes haptic input via an internal palpation-to-pain mapping model. To evaluate perceptual congruence between palpation and the corresponding auditory output, we conducted a study involving 7680 trials across 20 participants, where they evaluated pain intensity through sound. Results show that amplitude and pitch significantly influence agreement with the robot’s pain expressions, irrespective of pain sounds. Stronger palpation forces elicited stronger agreement, aligning with psychophysical patterns. The study revealed two key dimensions: pitch and amplitude are central to how people perceive pain sounds, with pitch being the most influential cue. These acoustic features shape how well the sound matches the applied force during palpation, impacting perceived realism. This approach lays the groundwork for high-fidelity robotic patients in clinical education and diagnostic simulation.
@@ 1 Introduction
Misdiagnosis is a critical global health issue [1], resulting in an estimated 795,000 cases of permanent harm or death annually in the United States alone [2]. The economic burden of diagnostic errors in the U.S. healthcare system is estimated to exceed \$100 billion per year [3]. These figures highlight the urgent need for effective training tools that improve diagnostic accuracy and reduce preventable harm. In this context, robotic patients or robopatients provide a high-fidelity, standardized medical training simulators for healthcare professionals. Unlike human actors, robopatients can consistently simulate rare or subtle conditions, allowing for repeated exposure and skill refinement.
@@
Palpation is a critical diagnostic skill where clinicians rely heavily on interpreting subtle pain responses. Synthetic pain generation—combining facial expressions, vocalizations, and haptic responses—is especially crucial in training scenarios such as palpation, where visual or auditory pain cues guide clinical judgment [4, 5]. By enabling controlled, multimodal feedback in response to tactile input, robopatients help practitioners develop better sensitivity to patients’ pain responses. This approach enhances diagnostic precision and supports the development of empathy-driven care. However, accurately recognizing and interpreting another person’s pain based on multimodal information remains underexplored. This gap arises from challenges in defining pain within diverse ontological frameworks [6, 7], as well as the limited understanding of pain’s underlying physiology [8].
@@
Pain is often communicated through facial expressions [9], which are influenced not only by physiological states but also by social and contextual factors [10, 11]. Despite its inherently interdisciplinary nature—spanning neuroscience, psychology, social contexts, and computer science—the study of pain has largely lacked comprehensive cross-disciplinary approaches. This limitation highlights the need for novel exploratory paradigms to better understand and simulate pain, particularly in controlled environments such as medical training systems.
@@
Figure 1: Multimodal pain synthesis for robopatient. The system comprises three main components: the user (ideally a physician), the robopatient, and the human, providing feedback to train the robopatient. The human patients under examination may exhibit various physiological abnormalities, psychological conditions, and individual characteristics, all of which influence the dimensionality of pain expression in response to palpation. The robopatient serves as an adaptive intermediary, learning appropriate pain responses from the user and aiming to replicate human pain expressions (including vocal pain sounds and facial pain expressions) realistically. During palpation, robot’s abdominal phantom captures user’s haptic input, which is recorded and to be processed through the robopatient’s internal palpation-pain mapping model. This model is derived from data collected from participants of diverse backgrounds. Our findings inform a universal method for mapping palpation to pain expressions, formulated as the “auditory-tactile congruence model”. Then the robot outputs corresponding pain sounds and facial expressions through MorphFace [12]. The user provides feedback on the robopatient’s generated pain responses, which was later refined by comparing palpation data with user feedback to map the relationship between palpation and the pain expressions it causes. This concept was adapted from [4] and modified to fit the new paradigm.
@@
Due to the risks associated with training on actual patients in nursing situations, simulation-based education (SBE) is used to train medical students. SBE enables a safe and controlled environment [13] for medical students to practice their physical examination including abdominal palpation. Social science rises visibly above the level of common sense in contributing to our understanding of communication in the patient-physician relationship [14]. Standardised patients (SPs) or professionally trained actors acting as patients, have been the highest fidelity simulators used in the medical field for decades [15]. SPs were capable of replicating facial expressions, vocal or audio expressions, both in the form of meaningful sentences or unstructured grunts, and variations of muscle stiffness during palpation training. Although this is effective, it is time intensive to train and maintain their skills. Hence we introduced robopatient [12], a robotic medical training simulator that helps ground the emotions associated with a patient and aid the diagnosis process. This further offers a platform for studying complex subjective psychophysiological phenomena such as pain by translating them into interpretable embodied expressions through robots.
@@
In this work we introduce a novel approach for generating synthetic multimodal pain expressions by integrating human feedback. This platform combines dynamic vocal pain sounds and facial expressions to simulate humanlike pain responses corresponding to palpation forces. This concept visualized in Figure 1, comprises a force-sensing abdominal phantom equipped with load cells and a controllable 3D hybrid face, MorphFace [12], capable of displaying facial and vocal pain expressions. An experimental study was conducted involving 7680 trials across 20 participants (10 males, 10 females) performing palpation tasks on robopatient across set target force levels. The real-time auditory and visual feedback was generated by the robot directly mapping the palpation to pain and allowed participants to rate the perceived intensity of pain sounds. Selected properties of pain expressions were changed to find the contribution of gender (male and female sounds) and acoustic features (amplitude and pitch) of pain sounds towards perceived pain during palpation. This systematic investigation into the relationship between properties of pain sounds and force levels offers a unique framework for exploring pain perception in medical training.
@@ 2 Related Work
Robots can study complex psychophysiological phenomena such as pain by simulating human-like responses through biologically inspired models. The Brain-inspired Robot Pain Spiking Neural Network (BRP-SNN) mimicing neural pathways involved in pain perception is an example for enabling robots to exhibit adaptive pain responses to physical stimuli [16]. Additionally, the care training assistant robot with 3D facial pain expressions, enhancing caregiver training by providing realistic, visual pain cues [17]. These robotic systems allow for controlled, repeatable studies of pain mechanisms and improve training by offering consistent, interpretable simulations of human pain responses.
Computational pain recognition has two sides: clinical and social and clinical aspect of pain includes pain detection and assessment of pain intensity [18]. Humans vocalise different types of pain sounds in different contexts such as vocal pain expressions during abdominal palpation and when someone touches a hot object. Hence it makes modelling the pain sounds or vocal pain expressions challenging [19, 20]. In addition, pain perception and expressions vary depending on human backgrounds, such as gender [21] and ethnicity [22]. When infants experience a higher intensity of pain, they produce more irregular cries [23] with higher amplitude [24, 25], larger fundamental frequency variation [26, 27], smaller amplitude variation, and longer duration [26, 28]. Hence the computational pain recognition must account for both clinical assessment and social context, as vocal pain expressions vary widely based on the situation, individual background (e.g., gender, ethnicity), and intensity of pain. This variability makes modeling pain sounds complex and context-dependent.
@@
Facial pain signals can sometimes operate independently of voice generation, influenced by features such as fundamental phonation frequencies and intonation [29]. However, sound has been found to be more informative than visual signals in expressing emotions [30], making it a vital modality for simulating pain in computational assistive systems. By incorporating multimodal signals, including synthetic vocal expressions that mimic human pain, robotic systems can foster humanlike empathy in contexts such as nursing and healthcare [31]. Despite its inherently interdisciplinary nature—spanning neuroscience, psychology, social contexts, and computer science—the study of pain and its synthetic representation has yet to fully take advantage of these connections. This highlights the need for innovative, cross-disciplinary approaches to better understand and simulate pain in training systems and assistive technologies.
@@
There exists a correlation between bio-signal parameters related to speech prosody and self-reported pain levels [32]. Therefore, nonlinguistic information or “how” we speak could sometimes be more important than the linguistic information or “what” we say. In addition, gestures could be the most accurate indicator of pain in an unimodal emotion recognition system, followed by speech and facial expressions [33]. Hence gestures could be accurately read by a doctor during the palpation while sound and facial expressions can be confusing. Robopatient simulates these two confusing features in order to assist doctors recognise pain demonstrated by face and nonlinguistic sound. For this, understanding human pain is crucial.
Multimodal information is more advantageous than unimodal in identifying pain from facial expressions, head posture, and physiological signals in videos[34]. There are different pain assessment scales used at present to decipher pain; the Adult Non-Verbal Pain Scale (NVPS) for instance takes facial expressions, movements, guarding, vital signs etc. in to account. Similarly, the Visual Analogue Scale (VAS) considers the amount of pain that a patient feels ranges across a continuum from none to an extreme amount of pain. Behavioral Pain Scale (BPS)- facial expressions, upper limbs, compliance. The Numeric Pain Rating Scale with the ventilator patient is asked to make three pain ratings, corresponding to current, best and worst pain experienced over the past 24 hours- none, mild, moderate to severe. As quantifying pain is not straight forward, we selected different palpation force levels that indicate different levels of pain.
@@
Due to the complexities involved in modeling pain sounds in humans, only a handful of studies have been reported. A recent study on pain analysed the acoustic characteristics which are responsible for variation in pain ratings in pain simulation by trained actors [35]. This study identified a range of voice fundamental frequency (F0, perceived as pitch), the amplitude of the vocalisation, the degree of periodicity of the vocalisation and the proportion of the signal displaying non-linear phenomena all increased with the level of simulated pain intensity. Hence increased the perceived pain by the listeners. Although these actors closely imitated infants and nonhuman mammal pain vocalization, pain studies related to adult humans are scarce.
@@
How robots recognise pain signals, such as humans’ painful expressions, and realise pain empathy for humans people tend to change their biophysical signals to improve the extent to which their minds are being “read”. Pain make sense in robots in two different forms: pain detection and pain generation. Our study is on pain generation for a robopatient which should cover diverse ethnicities, genders and ages. Elicited expressions which indicate expressions evoked by audiovisual media or real interpersonal interaction have been generated and projected on the robopatient. Gender differences in the experience and processing of emotion contribute to differences in the modulation and perception of pain due to differences in the modulation and perception of pain [36]. Hence we conducted our study with equal number of participants from each gender to keep gender balance and study gender-related trends as well. In addition there have been robotic platforms which studied concepts underlying physical pain such as SEATTLE [37] which resembles an artificial “pain nervous system”. However, the multidimensional nature of pain—extending beyond the nervous system—remains poorly understood.
@@
Considering the above aspects of palpation and pain expression, we tailored our study to gain insights from people to develop a robopatient which could replicate vocal and facial expressions related to pain during palpation. This could be used as a training simulator to aid doctors and medical students diagnose illnesses based on a patient’s vocal and facial expression of pain. Hence the goal of this research was to model the dynamic pain through voice and facial expression during palpation. We identified that there are no subsequent models which could map voice and the facial expressions simultaneously to pain during palpation. To develop such a model, we hypothesized that data-driven methods have the capability to model the relationship between the palpation behaviour, and vocal and facial pain expressions, with the help of reverse correlation and human user studies. To this end, we conducted a study where the participants could select appropriate pain expressions in terms of voice and facial expressions to represent different types of pain created by palpation. We used the force applied during the palpation as a parameter to demonstrate the level of pain a patient will experience during palpation. In the end, we present the haptic-auditory-visual scenario of a patient in our model during palpation.
@@ 3 Synthetic Pain Generation in Robopatient
Figure 2: Physical robot and the experiment procedure A) Robopatient setup: A scenario where the participant is palpating the robopatient is shown. B) Data flow within the robopatient setup during the experiment. The interaction between the robot and the user starts as the user palpates the abdominal phantom (Steps 1,2). The force the user applies has to reach a certain level indicated by a progress bar next to the robot’s face (Step 3). The progress bar is initially green and turns red when the required minimum amount of force is applied. Based on the palpation force, the robot generates facial expressions and sounds independently of each other (Step 4). Then the user listens to the pain sounds generated by the robot for the palpation force (Steps 5, 6) and provides feedback on whether they agree with the robot’s mapping of pain sounds to palpation.
@@ 3.1 Selection of Pain Sounds
For the experiment, three different pain sounds were selected per gender. This is to observe whether participants perceived certain pain sounds better than others. The pain sounds that were deemed appropriate were selected from the pilot trials. The pain sounds were selected from an online sound effects platform (Soundsnap). Male and female pain sounds were played to observe whether participants of a certain gender (male or female) may perceive the corresponding pain sounds better or worse. The audioread function in MATLAB 2024a (Mathworks, Inc.) was used to import audio files and individually obtain the amplitude and frequency (pitch) of the audio. The sound function in MATLAB was used to play the pain sounds.

### v2

@@ Abstract
Misdiagnosis can lead to delayed treatments and harm. Robopatients provide a controlled method for training and evaluating clinicians in rare, subtle, and complex cases, thereby reducing diagnostic errors. We present a study exploring auditory-tactile congruence for the synthesis of adaptive pain expressions in robopatients. In this work, the robopatient functions as an adaptive intermediary, capable of synthesizing vocal pain expressions in response to tactile stimuli generated during palpation. Using an abdominal phantom, robopatient captures and processes haptic input via an internal palpation-to-pain mapping model. To evaluate perceptual congruence between palpation and the corresponding auditory output, we conducted a study involving 7680 trials across 20 participants, who evaluated pain intensity through sound. Results show amplitude and pitch significantly influence agreement with the robot’s pain expressions, irrespective of pain sounds. Stronger palpation forces elicited stronger agreement, aligning with psychophysical patterns. The study revealed that pitch and amplitude are key in pain sound perception, with pitch being more influential. These acoustic features shape how well the sound matches the applied force during palpation, impacting perceived realism. This approach lays the groundwork for high-fidelity robopatients in clinical education and diagnostic simulation, serving as a platform to explore high-dimensional concepts such as pain through robots.
@@ 1 Introduction
Misdiagnosis is a critical global health issue [45], resulting in an estimated 795,000 cases of permanent harm or death annually in the United States alone [33]. The economic burden of diagnostic errors in the U.S. healthcare system is estimated to exceed \$100 billion per year [3]. These figures highlight the urgent need for effective training tools that improve diagnostic accuracy and reduce preventable harm. Furthermore, the UK government’s 10-year plan published in 2025, highlights the use of assistive robotic technologies and aims to reduce the need for hospital-based diagnostics [14]. In this context, robotic patients or robopatients provide a high-fidelity, standardized medical training simulators for healthcare professionals. Unlike human actors, robopatients can consistently simulate rare or subtle conditions, allowing for repeated exposure and skill refinement.
@@
Palpation is a critical diagnostic skill where clinicians rely heavily on interpreting subtle pain responses. Synthetic pain generation, combining facial expressions, vocalizations, and haptic responses, is especially crucial in training scenarios such as palpation, where visual or auditory pain cues guide clinical judgment [24, 39]. By enabling controlled, vocal pain feedback in response to tactile input, robopatients help practitioners develop better sensitivity to patients’ pain responses. This approach enhances diagnostic precision and supports the development of empathy-driven care. However, accurately recognizing and interpreting another person’s pain based on vocal pain information remains underexplored. This gap arises from challenges in defining pain within diverse ontological frameworks [42, 30], as well as the limited understanding of pain’s underlying physiology [15].
@@
Pain is often communicated through facial expressions [37], which are influenced not only by physiological states but also by social and contextual factors [51, 12]. Despite its inherently interdisciplinary nature, spanning neuroscience, psychology, social contexts, and computer science, the study of pain has largely lacked comprehensive cross-disciplinary approaches. This limitation highlights the need for novel exploratory paradigms to better understand and simulate pain, particularly in controlled environments such as medical training systems.
@@
Figure 1: Overview of the conceptual model for studying auditory-tactile congruence in robopatients. The system comprises three main components: the user (ideally a physician), the robopatient, and the human, providing feedback to train the robopatient. The human patients under examination may exhibit various physiological abnormalities, psychological conditions, and individual characteristics, all of which influence the dimensionality of pain expression in response to palpation. The robopatient serves as an adaptive intermediary, learning appropriate pain responses from the user and aiming to replicate human vocal pain expressions (including facial pain expressions for realism) realistically. During palpation, the robot’s abdominal phantom captures the user’s haptic input, which is recorded and processed through the robopatient’s internal palpation-pain mapping model. This conceptual model is derived from data collected from participants of diverse backgrounds during previous studies [24, 39, 46], and these studies inform a universal method for mapping palpation to pain expressions, formulated as the “auditory-tactile congruence model”. Then the robot outputs corresponding pain sounds and facial expressions through MorphFace [25]. The user provides feedback on the robopatient’s generated pain responses, which were later refined by comparing palpation data with user feedback to map the relationship between palpation and the pain expressions it causes. This concept was adapted from [24] and modified to fit the new paradigm.
@@
Due to the risks associated with training on actual patients in nursing situations, simulation-based education (SBE) is used to train medical students. SBE enables a safe and controlled environment [1] for medical students to practice their physical examination, including abdominal palpation. Social science rises visibly above the level of common sense in contributing to our understanding of communication in the patient-physician relationship [8]. Standardized patients (SPs) or professionally trained actors acting as patients, have been the highest fidelity simulators used in the medical field for decades [10]. SPs were capable of replicating facial expressions, vocal or audio expressions, both in the form of meaningful sentences or unstructured grunts, and variations of muscle stiffness during palpation training. Although this is effective, it is time-intensive to train and maintain their skills. Hence, we introduced robopatient [25], a medical robotic training simulator that helps ground the emotions associated with a patient and aid the diagnosis process. This further offers a platform for studying complex subjective psychophysiological phenomena, such as pain, by translating them into interpretable embodied expressions through robots.
@@
In this work, we introduce a novel approach for generating vocal pain expressions by integrating human feedback. This platform combines dynamic vocal pain sounds and facial expressions to simulate human-like pain responses corresponding to palpation forces. This concept, visualized in Figure 1, comprises a force-sensitive abdominal phantom equipped with load cells and a controllable 3D hybrid face, MorphFace [25], capable of displaying facial and vocal pain expressions. An experimental study was conducted involving 7680 trials across 20 participants (10 males, 10 females) performing palpation tasks on a robopatient across set target force levels. The real-time auditory and visual feedback was generated by the robot, directly mapping the palpation to pain, and allowed participants to rate the perceived intensity of pain sounds. Selected properties of pain expressions were changed to find the contribution of gender (male and female sounds) and acoustic features (amplitude and pitch) of pain sounds towards perceived pain during palpation. This systematic investigation into the relationship between properties of pain sounds and force levels offers a unique framework for exploring pain perception in medical training.
The rest of the paper is organized as follows: The prior work and its limitations are discussed in section 2. The implementation of the synthetic pain sounds, their selection, and the hardware system of the robopatient are discussed in section 3. The experiment methods and protocol used to conduct the study with human participants are described in section 4. The results of the study are discussed in section 5 with the data curation and analysis techniques used. The post-study questionnaires completed by the participants are also analyzed and discussed in this section. The limitations of the current study, its implications, and the future directions of this work are also mentioned in this section. Finally, we conclude in section 6 by discussing the key findings of this work and its contributions that could lead to the development of robotic patients for medical training.
@@ 2 Related Work
Robots can be useful for studying complex psychophysiological phenomena, such as pain, by simulating human-like responses through biologically inspired models. The Brain-inspired Robot Pain Spiking Neural Network (BRP-SNN) that mimics the neural pathways involved in pain perception is an example of enabling robots to exhibit adaptive pain responses to physical stimuli [16]. Additionally, a care training assistant robot with 3D facial pain expressions can enhance caregiver training by providing realistic visual pain cues [27]. These robotic systems allow for controlled, repeatable studies of pain mechanisms and improve training by offering consistent, interpretable simulations of human pain responses.
@@
Computational pain recognition has two sides: clinical and social. The clinical aspect of pain includes the detection and assessment of the intensity of pain [6]. Humans vocalize different types of pain sounds in different contexts, such as vocal pain expressions during abdominal palpation and when someone touches a hot object. Hence, modeling pain sounds or vocal pain expressions is challenging [26, 19]. In addition, pain perception and expressions vary depending on human backgrounds, such as gender [43] and ethnicity [40]. When infants experience a higher intensity of pain, they produce more irregular cries [48] with higher amplitude [28, 17], larger fundamental frequency variation [36, 23], smaller amplitude variation, and for longer duration [36, 20]. Hence, the computational pain recognition must account for both clinical assessment and social context, as vocal pain expressions vary widely based on the situation, individual background (e.g., gender, ethnicity), and intensity of pain. This variability makes modeling pain sounds complex and context-dependent. Facial pain signals can sometimes operate independently of voice generation, influenced by features such as fundamental phonation frequencies and intonation [5]. However, sound is more informative than visual signals in expressing emotions [49], making it a vital modality for simulating pain in computational assistive systems. By incorporating multimodal signals, including synthetic vocal expressions that mimic human pain, robotic systems can foster human-like empathy in contexts such as nursing and healthcare [35]. Despite its inherently interdisciplinary nature spanning neuroscience, psychology, social contexts, and computer science, the study of pain and its synthetic representation has yet to fully take advantage of these connections. This highlights the need for innovative, cross-disciplinary approaches to better understand and simulate pain in training systems and assistive technologies.
@@
There exists a correlation between bio-signal parameters related to speech prosody and self-reported pain levels [34]. Therefore, non-linguistic information or “how” we speak could sometimes be more important than the linguistic information or “what” we say. In addition, gestures could be the most accurate indicator of pain in an unimodal emotion recognition system, followed by speech and facial expressions [7]. Hence, gestures could be accurately read by a doctor during the palpation, while sound and facial expressions can be confusing. Robopatients, on the other hand can simulate these two confusing features in order to assist doctors in recognizing pain demonstrated by facial expressions and non-linguistic sounds. For this, understanding human pain is crucial.
@@
Due to the complexities involved in modeling pain sounds in humans, only a handful of studies have been reported. A recent study on pain analyzed the acoustic characteristics that are responsible for variation in pain ratings in pain simulation by trained actors [41]. This study identified that a range of voice fundamental frequency (F0, perceived as pitch), the amplitude of the vocalization, the degree of periodicity of the vocalization, and the proportion of the signal displaying non-linear phenomena all increased with the level of simulated pain intensity. Thus increasing the perceived pain by the listeners. Although these actors closely imitate infants and nonhuman mammals’ pain vocalization, pain studies related to adult humans are scarce.
@@
When robots recognize human pain signals, such as painful facial expressions, and display pain-related empathy, people often adjust their own biophysical signals to make their internal state more easily ‘read’. Pain make sense in robots in two different forms: pain detection and pain generation. Our study is on pain generation for a robopatient, which should cover diverse ethnicities, genders, and ages. Elicited expressions, which indicate expressions evoked by audiovisual media or real interpersonal interaction, have been generated and projected on the robopatient. Gender differences in the experience and processing of emotion contribute to differences in the modulation and perception of pain due to differences in the modulation and perception of pain [44]. In addition, there have been robotic platforms that studied concepts underlying physical pain, such as SEATTLE [2], which resembles an artificial “pain nervous system”. However, the multidimensional nature of pain, which extends beyond the nervous system, remains poorly understood.
@@
Considering the above aspects of palpation and pain expression, we tailored our study to gain insights from people to develop a robopatient that could replicate vocal and facial expressions related to pain during palpation. We identified that there are no subsequent models that could map voice and facial expressions simultaneously to pain during palpation. To develop such a model, we hypothesized that data-driven methods can model the relationship between the palpation behaviour and vocal expressions, with the help of reverse correlation and user studies. A similar data-driven approach was used in prior work, which was based on audio-tactile cross-modal congruence for surface texture classification [29]. To this end, we conducted a study where the participants could select appropriate pain expressions in terms of voice and facial expressions to represent different types of pain created by palpation. We used the force applied during the palpation as a parameter to demonstrate the level of pain a patient will experience during palpation. In the end, we present the haptic-auditory-visual scenario of a patient in our model during palpation.
@@ 3 Synthetic Pain Generation in Robopatient
Figure 2: Physical robot and the experiment procedure A) Robopatient setup. A scenario where the participant is palpating the robopatient is shown. B) Data flow within the robopatient setup during the experiment. The interaction between the robot and the user starts as the user palpates the abdominal phantom (Steps 1, 2). The force the user applies has to reach a certain level indicated by a progress bar next to the robot’s face (Step 3). The progress bar is initially green and turns red when the required minimum amount of force is applied. Based on the palpation force, the robot generates facial expressions and sounds independently of each other (Step 4). Then the user listens to the pain sounds generated by the robot for the palpation force (Steps 5, 6) and provides feedback on whether they agree with the robot’s mapping of pain sounds to palpation.
@@ 3.1 Selection of Pain Sounds
For the experiment, three different pain sounds were selected per gender. This is to observe whether participants perceived certain pain sounds better than others. The pain sounds that were deemed appropriate were selected from the pilot trials. The pain sounds were selected from an online sound effects platform (Soundsnap). Male and female pain sounds were played to observe whether participants of a certain gender (male or female) may perceive the corresponding pain sounds better or worse. The audioread function in MATLAB 2024a (Mathworks, Inc.) was used to import audio files and individually obtain the amplitude and frequency (pitch) of the audio. The sound function in MATLAB was used to play the pain sounds. The male and female sounds selected for the experiment mentioned in section 4 are provided here with their designated pain sound numbers (i.e., 1, 2, and 3) appended to the end of their filenames.